# score-based generative model SNR signal-to-noise ratio

### v2

@@ I Introduction
While the task of neural speech enhancement (SE) has traditionally been considered a discriminative task, works from recent years, in particular SGMSE+ [richter2023speech], have shown that treating SE as a generative task instead can have several advantages, particularly in perceived speech quality and model generalization. Furthermore, other speech processing tasks such as bandwidth extension (BWE) or Mel vocoding are naturally well-suited to be treated as generative tasks, since significant amounts of information are missing from the input signal and must be regenerated given only the remaining information [lemercier2023analysing]. Other tasks which have been successfully tackled from this point of view include neural codec post-filtering [wu2024scoredec, welker2025flowdec], dereverberation [richter2023speech, lemercier2023analysing, lemercier2023storm], short-time Fourier transform (STFT) phase retrieval (PR) [peer2023diffphase], and binaural speech synthesis [liang2025binauralflow].
@@
However, a key downside to such generative methods is that they are computationally intensive due to the involved numerical solvers evaluating a large deep neural network (DNN) multiple times. This seems to preclude the use of these methods for many real-time scenarios [lay2025diffusion]. However, we show here that real-time and multi-step inference are not necessarily at odds with each other and can be realized on available consumer hardware, given that one uses appropriately buffered inference schemes and matching network architectures.
@@
While many prior works describe their methods as real-time [liang2025binauralflow, hsieh2025towards, lu2025realtime], this is often either (1) not supported through timings on real hardware, making it unclear whether a streaming implementation is realistically attainable, or (2) measured based on the real-time factor (RTF) estimated using offline processing, i.e., using an input utterance duration longer than a single frame, and dividing by the processing time the model takes for the full utterance. This neglects that, in offline processing, the inference process can make full use of parallelism across the time dimension of the input signal, as well as CUDA kernels which are typically optimized to process single large tensors as quickly as possible rather than many smaller tensors one-by-one. Such reported offline RTFs may therefore severely underestimate the actual RTF for streaming inference, making it unclear which methods are practically real-time capable in a streaming setting.
@@
With this work, we aim to close this gap and make modern generative streaming-capable methods available to the research community. We propose Stream.FM, a real-time capable streaming generative model based on flow matching [lipman2023flow] that can solve various speech restoration problems with a total latency below 50 ms, bringing the high-quality capabilities of diffusion-based generative models to real-time speech processing. We expand upon our previous conference submission [welker2025realtime] for real-time Mel vocoding in the following ways: (1) we investigate five additional speech processing tasks beyond Mel vocoding: speech enhancement, dereverberation, bandwidth extension, codec post-filtering, and STFT phase retrieval; (2) for SE, we make use of a predictive-generative model combination as proposed in StoRM [lemercier2023storm]; (3) we newly propose and investigate the use of learned Runge-Kutta ordinary differential equation (ODE) solvers to boost model output quality under a fixed computational budget; (4) we explore the use of model weight compression towards a flexible and favorable choice along the computation/quality tradeoff.
The predictive-generative SE model detailed here was presented in a real-time demonstration at the ITG Conference on Speech Communication [welker2025itgdemo]. Code [^1] https://github.com/sp-uhh/streamfm, published after acceptance. and audio examples are available online [^2] https://sp-uhh.github.io/streamfm_examples.
@@ II-A Diffusion- and flow-based speech processing
Diffusion-based speech enhancement was originally introduced in [lu2022conditional, welker2022speech] and first achieved state-of-the-art quality with SGMSE+ [richter2023speech]. Song et al. [song2021scorebased] originally proposed to define diffusion models for generative modeling via time-continuous forward stochastic differential equations that model a continuous mapping from data to noise. Each forward SDE has a corresponding reverse SDE resulting directly from mathematical theory [song2021scorebased], which can be used to map from tractable noise samples to newly generated data. The reverse SDE involves the score function $\nabla_{x}\log(p(x))$ of the data distribution $p(x)$, which is intractable in general but can be learned by a neural network called a score model [song2021scorebased].
For speech enhancement, SGMSE [welker2022speech] and SGMSE+ [richter2023speech] propose specific modified SDEs, modeling speech corruption by interpolating between clean speech $s$ and corrupted speech $y$ and incrementally adding Gaussian white noise. To produce enhanced speech, one draws a Gaussian white noise sample, adds it to the corrupted speech, and then numerically solves the reverse SDE starting from this point. We refer the reader to [richter2023speech] for the full detailed description of the training and inference.
Later works introduce flow matching (FM) [lipman2023flow], which is closely related to diffusion models but takes a perspective based on ODEs rather than SDEs. The idea is to learn a model to transport samples from a tractable distribution $q_{0}(X_{0})$, e.g., a multivariate Gaussian, to an intractable data distribution $q_{1}(X_{1})=p_{\rm{data}}$ by solving the ODE
@@
starting from a random sample $X_{0}\sim q_{0}$. $\phi:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is called the flow with the associated time-dependent vector field $u:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$, generating a probability density path $p_{\tau}:\mathbb{R}^{n}\to\mathbb{R}_{+}$, i.e., a probability density function that depends on an artificial time coordinate $\tau\in[0,1]$ with $p_{0}=q_{0}$ and $p_{1}=q_{1}$. One can learn a neural network $v_{\theta}$ called a flow model to approximate $u(\tau,\cdot)$ with the conditional flow matching loss [lipman2023flow, Eq. 9]:
@@
where $\tau\sim\mathcal{U}(0,1),X\sim q_{1}$ is clean data sampled from a training corpus, $\mathbb{E}_{a}[b(a)]=\int_{-\infty}^{\infty}b(a)p(a)\mathrm{d}a$ denotes the expectation of $b(a)$ with respect to the distribution $p(a)$ of $a$, and $\theta$ denotes the set of DNN parameters. Note that, during training, the expectation is approximated by an empirical average over the training data. The objective Eq. 2, where $u$ is conditional on the clean $X$, has the same gradients as an intractable objective where $u$ is unconditional [lipman2023flow, Eq. 5], and results in the correct probability path $p_{\tau}(X_{\tau})$ and flow field $u(\tau,X_{\tau})$ [lipman2023flow, Sec. 3.1, 3.2]. FlowDec [welker2025flowdec] showed that, similar to SGMSE [welker2022speech, richter2023speech], FM can be modified to interpolate between the distributions of clean and corrupted data, enabling the use of FM for generative signal enhancement. Concretely, we choose the following path, which linearly interpolates from a clean sample $S$ to a corrupted sample $Y$ and adds increasing amounts of Gaussian noise:
$p_{\tau}(X_{\tau}|S,Y):=\mathcal{N}(X_{\tau};(1-\tau)Y+\tau(S-Y),\Sigma_{\tau})$ (3)
where $\Sigma_{\tau}=((1-\tau)\Sigma_{y}+\tau\Sigma_{\text{min}})^{2}$ and $\Sigma_{y},\Sigma_{\text{min}}$ denote covariance matrices which we assume to be scalar or diagonal. We denote a scalar covariance by $\sigma_{y}^{2}$, i.e., $\Sigma_{y}=\sigma_{y}^{2}I$ with the identity matrix $I$. The flow model $v_{\theta}$ can be trained via the joint flow matching loss [welker2025flowdec]:
@@
where $(S,Y)$ are paired data of clean and corrupted audio sampled from a training corpus. Note that we let each set of $X_{1},X_{0}$ and $X_{\tau}$ share a single Gaussian noise sample $\varepsilon\sim\mathcal{N}(0,I)$, as a simple form of training variance reduction through minibatch coupling [pooladian2023multisample]. While [welker2025flowdec] did not use $\Sigma_{\text{min}}$, we reintroduce it here similar to [lipman2023flow] as we found this to slightly increase training stability.
@@ II-B Latency definitions
We define the algorithmic latency $\ell_{\text{alg}}$ as the attainable latency on infinitely fast hardware, and the overall latency of a system as $\ell_{\text{tot}}=\ell_{\text{alg}}+\ell_{\text{proc}}$, where $\ell_{\text{proc}}$ is the processing time per frame. For a frame-causal STFT-based method with causal (right-hand) window alignment such as Stream.FM, the algorithmic latency $\ell_{\text{alg}}$ is the synthesis window length $W_{\text{syn}}$ minus one sample divided by the sampling rate $f_{s}$, i.e., $\ell_{\text{alg}}=\frac{W_{\text{syn}}-1}{f_{s}}$ [wood2019unsupervised, wang2023verylow]. In typical frame-by-frame processing implementations, $\ell_{\text{proc}}$ is effectively given exactly via the synthesis frame hop $H_{\text{syn}}$, due to the synthesis side waiting for each synthesis frame hop to be complete before producing audio samples. This allows the processing model up to $\ell_{\text{proc}}\leq\frac{H_{\text{syn}}}{f_{s}}$ of processing time. We use the same configurations for analysis and synthesis here, hence the analysis window $W_{\text{ana}}=W_{\text{syn}}$ and hop $H_{\text{ana}}=H_{\text{syn}}$, so $\ell_{\text{tot}}=\frac{W_{\text{ana}}-1+H_{\text{ana}}}{f_{s}}$ for our models.
@@ II-C Explicit Runge-Kutta ODE solvers
To solve ODEs such as Eq. 1 numerically, a simple method is the Euler method [hairer1993solving]. It starts from $X_{0}$ and $\tau=0$, discretizes the time interval $\tau\in[0,1]$ into uniformly spaced points, and iterates the following equation for $N$ iterations until $\tau=1$ using $\Delta\tau=\frac{1}{N}$:
@@
where we inserted the learned flow model $v_{\theta}$ for $u$ in the approximate equality. The Euler solver thus requires $\text{NFE}=N$ calls of the DNN $v_{\theta}$, but has a relatively large approximation error for a given NFE budget compared to more flexible solvers [hairer1993solving]. For high-quality real-time inference, we want our solvers to have a fixed small NFE and attain high quality under this budget. To this end, we use explicit Runge-Kutta solvers [hairer1993solving], which can be parameterized by a set of coefficients $\mathbf{A}\in\mathbb{R}^{r\times r},\mathbf{b}\in\mathbb{R}^{r},\mathbf{c}\in\mathbb{R}^{r}$ [butcher1964highorder] where $r$ is the NFE per step, and $\mathbf{A}$ is a strictly lower triangular matrix so the solver is explicit [hairer1993solving]. While these coefficients are usually predefined, in Section III-E, we propose to learn task-optimized $\{\mathbf{A},\mathbf{b},\mathbf{c}\}$ from data. The iterated equation of such a solver is:
@@
where $a_{ij}=\mathbf{A}[i,j]$. We set $N=1$ and $\Delta\tau=1$ in the following so that we perform a single evaluation of Eqs. 6, 7 and 8 at $\tau=0$, stepping directly to $\tau=1$. This yields $\text{NFE}=r$, with each $G_{i}\in\{G_{1},\ldots,G_{r}\}$ contributing one evaluation of $v_{\theta}$.
II-D Related work
In 2020, Défossez et al. [defossez2020realtime] proposed a predictive waveform-domain model for causal real-time SE, which we refer to as DEMUCS here. The recent work aTENNuate [pei2025optimized] is a predictive state-space model for real-time SE. DEMUCS and aTTENuate operate at a similar, slightly larger algorithmic latency than Stream.FM, but no streaming implementation of aTTd ENuate has been published, hence we only compare against an offline variant of it. Dmitrieva and Kaledin [dmitrieva2025hifi] recently proposed HiFi-Stream for streaming SE, which uses a block-wise inference scheme at the cost of possible block-edge discontinuities and a higher latency.
Richter et al. [richter2024causal] first proposed the use of causal DNNs for diffusion-based SE and achieved convincing results, but did not target real-time capability on real hardware, and used heavy down- and upsampling along time which leads to large latencies [hedegaard2022cins, lay2025diffusionjpre].
Lay et al. [lay2025diffusion] proposed the Diffusion Buffer (DB), which realizes a real-time streaming SE diffusion model on consumer hardware. DB couples the diffusion time $\tau$ to physical time $t$ in a fixed-size frame buffer. The buffer is progressively denoised frame-by-frame using one DNN call each time, and each time the oldest—then fully denoised—frame in the buffer is output. This realizes a fixed algorithmic latency equal to the total buffer duration. A key downside of DB is that the algorithmic latency, while sub-second, is still relatively large: a buffer of fewer than 20 frames did not yield satisfying quality, resulting in an effective minimum algorithmic latency of 320 ms in their configuration and around 180 ms in a follow-up preprint [lay2025diffusionjpre].
Liang et al. [liang2025binauralflow] describe a similar approach to streaming FM as detailed here, but only investigate monaural-to-binaural speech upmixing whereas we treat six different speech tasks. The authors determine RTFs on a 0.683-second snippet which may underestimate streaming RTF significantly, see Section V-F. To ensure reproducibility and enable future developments, we provide a more extensive description of the buffered multi-step inference scheme, see Section III-A, as well as a public code repository.
A concurrent work [hsieh2025towards] investigates streaming FM models for speech restoration, similarly using a U-Net architecture without time-wise downsampling as proposed here. The authors report an algorithmic latency of 20 ms, but do not provide real-hardware timings, making it unclear whether a streaming model can be practically realized.

### v3

@@
VB-DMD
VoiceBank-DEMAND
EWv2
EARS-WHAM v2
LRK
Learned Runge-Kutta
DB
Diffusion Buffer
@@ I Introduction
While the task of neural speech enhancement (SE) has traditionally been considered a predictive task, works from recent years, in particular SGMSE+ [47], have shown that treating SE as a generative task instead can have several advantages, particularly in perceived speech quality and model generalization. More broadly, non-additive speech restoration tasks such as bandwidth extension (BWE), dereverberation or short-time Fourier transform (STFT) phase retrieval (PR) are naturally well-suited for generative methods, in particular when a significant amount of information is missing from the input signal and must be regenerated given only the remaining information [30, 40]. Prior work [30] empirically supports this idea, showing clear advantages of modern generative methods over predictive methods for such problems. Other tasks which have been successfully tackled with diffusion-based generative methods include neural codec post-filtering [64, 60], Mel vocoding [61], and binaural speech synthesis [33].
@@
However, a key downside to such generative methods is that they are computationally intensive due to the involved numerical solvers evaluating a large deep neural network (DNN) multiple times. This seems to preclude the use of these methods for many real-time scenarios [27]. However, we show here that real-time and multi-step inference are not necessarily at odds with each other and can be realized on available consumer hardware, given that one uses appropriately buffered inference schemes and matching network architectures.
@@
While many prior works describe their methods as real-time [33, 17, 36], this is often either (a) not supported through timings on real hardware, making it unclear whether a streaming implementation is realistically attainable, or (b) measured based on the real-time factor (RTF) estimated using offline processing, i.e., using an input utterance duration longer than a single frame, and dividing by the processing time the model takes for the full utterance. This neglects that, in offline processing, the inference process can make full use of parallelism across the time dimension of the input signal, as well as CUDA kernels which are typically optimized to process single large tensors as quickly as possible rather than many smaller tensors one-by-one. Such reported offline RTFs may therefore severely underestimate the actual RTF for streaming inference, making it unclear which methods are practically real-time capable in a streaming setting.
With this work, we aim to close this gap and make modern generative streaming-capable methods available to the research community. We propose Stream.FM, a real-time capable streaming generative model based on flow matching [34] that can solve various speech restoration problems with a total latency below 50 ms, bringing the high-quality capabilities of diffusion-based generative models to real-time speech processing. We extend our previous conference publication on real-time streaming Mel vocoding [61], which (1) proposes a frame-wise causal DNN and iterative inference scheme for real-time flow matching inference, also used in this work, (2) realizes streaming inference with a total latency of 48 ms on consumer hardware, and (3) outperforms established non-causal baselines including HiFi-GAN [22].
@@
We expand upon our previous conference publication [61] in the following ways: (4) we demonstrate wide applicability of our model for general speech restoration with five additional restoration tasks beyond Mel vocoding: speech enhancement, dereverberation, bandwidth extension, codec post-filtering, and STFT phase retrieval; (5) we newly propose and investigate the use of learned Runge-Kutta ordinary differential equation (ODE) solvers for flow matching to boost output quality under a fixed computational budget; (6) we provide a comprehensive comparison of attainable streaming latencies of our model and baselines on consumer hardware; (7) for SE, we design a joint predictive-generative model inspired by StoRM [31]; (8) we explore the use of model weight compression [13] towards a flexible and favorable choice along the computation/quality tradeoff.
A full-band variant of the models detailed here, which jointly performs SE and BWE, has been accepted as a Show-and-Tell demo at ICASSP 2026 [59]. Code [^1] https://github.com/sp-uhh/streamfm, published after acceptance. and audio examples are available online [^2] https://sp-uhh.github.io/streamfm_examples.
I-A Related work
In 2020, Défossez et al. [9] proposed a predictive waveform-domain model for causal real-time SE, which we refer to as DEMUCS here. The recent work aTENNuate [42] is a predictive state-space model for real-time SE. DEMUCS and aTTENuate operate at a similar, slightly larger algorithmic latency than Stream.FM, but no streaming implementation of aTTENuate has been published. Dmitrieva and Kaledin [10] recently proposed HiFi-Stream for streaming SE, which uses a block-wise inference scheme at the cost of possible block-edge discontinuities and a higher latency. MambAttention [25] is a recently proposed model for generalizable speech enhancement, based on a combination of modern state-space models and attention layers, but is not targeted towards streaming inference and uses a non-causal configuration.
Richter et al. [48] first proposed the use of causal DNNs for diffusion-based SE and achieved convincing results, but did not target real-time capability on real hardware, and used heavy down- and upsampling along time which leads to large latencies [16, 28].
Lay et al. [27] proposed the Diffusion Buffer (DB), which realizes a real-time streaming SE diffusion model on consumer hardware. DB couples the diffusion time $\tau$ to physical time $t$ in a fixed-size frame buffer. The buffer is progressively denoised frame-by-frame using one DNN call each time, and each time the oldest—then fully denoised—frame in the buffer is output. This realizes a fixed algorithmic latency equal to the total buffer duration, which studies indicate should not be reduced below 180 ms [28].
Liang et al. [33] describe a similar approach to streaming flow matching (FM) as detailed here, but only investigate monaural-to-binaural speech upmixing whereas we treat six different speech tasks. The authors determine RTFs on a 0.683-second snippet which may underestimate streaming RTF significantly, see Section V-F. To ensure reproducibility and enable future developments, we provide a more extensive description of the buffered multi-step inference scheme, see Section III-A, as well as a public code repository.
AnyEnhance [68] is a generative speech restoration model using mask-based generative modeling of discrete speech tokens, and addresses various restoration tasks including denoising, dereverberation and bandwidth extension with a single model at a sampling rate of 44.1 kHz. While [68] shows impressive performance, it uses proprietary noise and RIR data which hinders reproducibility, and only a significantly smaller and less powerful model variant [^3] https://github.com/viewfinder-annn/AnyEnhance-v1 has been published with code and checkpoints as part of a challenge [69]. AnyEnhance performs enhancement in a highly compressed token space, rather than an uncompressed continuous space like our method. This can aid efficiency, but may also degrade the signal-level similarity to the clean target speech.
A concurrent work [17] investigates streaming FM models for speech restoration, similarly using a U-Net architecture without time-wise downsampling as proposed here. The authors report an algorithmic latency of 20 ms, but do not provide real-hardware timings, making it unclear whether a streaming model can be practically realized.
Note that all models described above are (a) not streaming-capable with low latency and/or (b) have only been demonstrated for speech enhancement – i.e., background noise removal – and not for general speech restoration tasks. This shows that a gap exists in the literature regarding streaming-capable models for general speech restoration, which we aim to close with this work.
@@ II-A Diffusion- and flow-based speech processing
Diffusion-based speech enhancement was originally introduced in [37, 62] and first achieved state-of-the-art quality with SGMSE+ [47]. Song et al. [54] originally proposed to define diffusion models for generative modeling via time-continuous forward stochastic differential equations that model a continuous mapping from data to noise. Each forward SDE has a corresponding reverse SDE resulting directly from mathematical theory [54], which can be used to map from tractable noise samples to newly generated data. The reverse SDE involves the score function $\nabla_{x}\log(p(x))$ of the data distribution $p(x)$, which is intractable in general but can be learned by a neural network called a score model [54].
For speech enhancement, SGMSE [62] and SGMSE+ [47] propose specific modified SDEs, modeling speech corruption by interpolating between clean speech $s$ and corrupted speech $y$ and incrementally adding Gaussian white noise. To produce enhanced speech, one draws a Gaussian white noise sample, adds it to the corrupted speech, and then numerically solves the reverse SDE starting from this point. We refer the reader to [47] for the full detailed description of the training and inference.
Later works introduce flow matching (FM) [34], which is closely related to diffusion models but takes a perspective based on ODEs rather than SDEs. The idea is to learn a model to transport samples from a tractable distribution $q_{0}(X_{0})$, e.g., a multivariate Gaussian, to an intractable data distribution $q_{1}(X_{1})=p_{\rm{data}}$ by solving the ODE
@@
starting from a random sample $X_{0}\sim q_{0}$. $\phi:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is called the flow with the associated time-dependent vector field $u:[0,1]\times\mathbb{R}^{n}\to\mathbb{R}^{n}$, generating a probability density path $p_{\tau}:\mathbb{R}^{n}\to\mathbb{R}_{+}$, i.e., a probability density function that depends on an artificial time coordinate $\tau\in[0,1]$ with $p_{0}=q_{0}$ and $p_{1}=q_{1}$. One can learn a neural network $v_{\theta}$ called a flow model to approximate $u(\tau,\cdot)$ with the conditional flow matching loss [34, Eq. 9]:
@@
where $\tau\sim\mathcal{U}(0,1),X\sim q_{1}$ is clean data sampled from a training corpus, $\mathbb{E}_{a}[b(a)]=\int_{-\infty}^{\infty}b(a)p(a)\mathrm{d}a$ denotes the expectation of $b(a)$ with respect to the distribution $p(a)$ of $a$, and $\theta$ denotes the set of DNN parameters. Note that, during training, the expectation is approximated by an empirical average over the training data. The objective Eq. 2, where $u$ is conditional on the clean $X$, has the same gradients as an intractable objective where $u$ is unconditional [34, Eq. 5], and results in the correct probability path $p_{\tau}(X_{\tau})$ and flow field $u(\tau,X_{\tau})$ [34, Sec. 3.1, 3.2]. FlowDec [60] showed that, similar to SGMSE [62, 47], FM can be modified to interpolate between the distributions of clean and corrupted data, enabling the use of FM for generative signal enhancement. Concretely, we choose the following path, which linearly interpolates from a clean sample $S$ to a corrupted sample $Y$ and adds increasing amounts of Gaussian noise:
$p_{\tau}(X_{\tau}|S,Y):=\mathcal{N}(X_{\tau};(1-\tau)Y+\tau S,\Sigma_{\tau})$ (3)
where $\Sigma_{\tau}=((1-\tau)\Sigma_{y}+\tau\Sigma_{\text{min}})^{2}$ and $\Sigma_{y},\Sigma_{\text{min}}$ denote covariance matrices which we assume to be scalar or diagonal. We denote a scalar covariance by $\sigma_{y}^{2}$, i.e., $\Sigma_{y}=\sigma_{y}^{2}I$ with the identity matrix $I$. The flow model $v_{\theta}$ can be trained via the joint flow matching loss [60]:
@@
where $(S,Y)$ are paired data of clean and corrupted audio sampled from a training corpus. Note that we let each set of $X_{1},X_{0}$ and $X_{\tau}$ share a single Gaussian noise sample $\varepsilon\sim\mathcal{N}(0,I)$, as a simple form of training variance reduction through minibatch coupling [43]. While [60] did not use $\Sigma_{\text{min}}$, we reintroduce it here similar to [34] as we found this to slightly increase training stability.
@@ II-B Latency definitions
We define the algorithmic latency $\ell_{\text{alg}}$ as the attainable latency on infinitely fast hardware, and the overall latency of a system as $\ell_{\text{tot}}=\ell_{\text{alg}}+\ell_{\text{proc}}$, where $\ell_{\text{proc}}$ is the processing time per frame. For a frame-causal STFT-based method with causal (right-hand) window alignment such as Stream.FM, the algorithmic latency $\ell_{\text{alg}}$ is the synthesis window length $W_{\text{syn}}$ minus one sample divided by the sampling rate $f_{s}$, i.e., $\ell_{\text{alg}}=\frac{W_{\text{syn}}-1}{f_{s}}$ [63, 58]. In typical frame-by-frame processing implementations, $\ell_{\text{proc}}$ is effectively given exactly via the synthesis frame hop $H_{\text{syn}}$, due to the synthesis side waiting for each synthesis frame hop to be complete before producing audio samples. This allows the processing model up to $\ell_{\text{proc}}\leq\frac{H_{\text{syn}}}{f_{s}}$ of processing time. We use the same configurations for analysis and synthesis here, hence the analysis window $W_{\text{ana}}=W_{\text{syn}}$ and hop $H_{\text{ana}}=H_{\text{syn}}$, so $\ell_{\text{tot}}=\frac{W_{\text{ana}}-1+H_{\text{ana}}}{f_{s}}$ for our models. We report both $\ell_{\text{alg}}$ and $\ell_{\text{tot}}$ for all methods, but note that to attain $\ell_{\text{tot}}$ on concrete hardware the streaming RTF must be reliably below 1, see Section V-F.
@@ II-C Explicit Runge-Kutta ODE solvers
To solve ODEs such as Eq. 1 numerically, a simple method is the Euler method [15]. It starts from $X_{0}$ and $\tau=0$, discretizes the time interval $\tau\in[0,1]$ into uniformly spaced points, and iterates the following equation for $N$ iterations until $\tau=1$ using $\Delta\tau=\frac{1}{N}$:
@@
where we inserted the learned flow model $v_{\theta}$ for $u$ in the approximate equality. The Euler solver thus requires $\text{NFE}=N$ calls of the DNN $v_{\theta}$, but has a relatively large approximation error for a given NFE budget compared to more flexible solvers [15]. For high-quality real-time inference, we want our solvers to have a fixed small NFE and attain high quality under this budget. To this end, we use explicit Runge-Kutta solvers [15], which can be parameterized by a set of coefficients $\mathbf{A}\in\mathbb{R}^{r\times r},\mathbf{b}\in\mathbb{R}^{r},\mathbf{c}\in\mathbb{R}^{r}$ [4] where $r$ is the NFE per step, and $\mathbf{A}$ is a strictly lower triangular matrix so the solver is explicit [15]. While these coefficients are usually predefined, in Section III-E, we propose to learn task-optimized $(\mathbf{A},\mathbf{b},\mathbf{c})$ from data. The iterated equation of such a solver is:
@@
where $a_{ij}=\mathbf{A}[i,j]$. We set $N=1$ and $\Delta\tau=1$ so that a single evaluation of Eqs. 6, 7 and 8 is performed, stepping directly to $\tau=1$. This guarantees $\text{NFE}=r$, with each $G_{i}$ contributing one evaluation of $v_{\theta}$.